Some systems are under maintenance.

ETH-Services Status

Status of AS200462

  • Network

    Last updated 2025-09-26 20:44:59

    Operational
  • Power

    Last updated 2025-09-18 17:23:01

    Operational
  • Cooling

    Last updated 2025-09-18 17:23:01

    Operational
  • AS174 - Cogent

    Last updated 2025-10-21 18:38:43

    Operational
  • AS1299 - Arelion

    Last updated 2025-11-10 12:30:53

    Under maintenance
  • AS2914 - NTT

    Last updated 2025-10-28 21:20:12

    Operational
  • AS5405 - Inter.Link

    Last updated 2025-11-10 12:30:55

    Operational
  • AS5511 - Orange

    Last updated 2025-11-11 22:18:37

    Operational
  • AS6453 - TATA

    Last updated 2025-11-11 22:18:38

    Operational
  • AS6762 - Telecom Italia Sparkle

    Last updated 2025-11-11 20:13:42

    Operational
  • AS6830 - Liberty Global

    Last updated 2025-11-11 22:18:39

    Operational
  • AS9002 - RETN

    Last updated 2025-11-11 22:33:49

    Under maintenance
  • AS34549 - Meerfarbig

    Last updated 2025-11-11 22:18:39

    Operational
  • Cloud

    Last updated 2025-09-18 17:23:01

    Operational
  • DNS

    Last updated 2025-10-23 12:38:37

    Operational
  • DDoS-Filter

    Last updated 2025-09-18 17:23:01

    Operational
  • E-Mail

    Last updated 2025-09-18 17:23:01

    Operational
  • VServer

    Last updated 2025-09-18 17:23:01

    Operational
  • VServer - Backups

    Last updated 2025-09-18 17:23:01

    Operational
  • Web

    Last updated 2025-09-18 17:23:01

    Operational

Past Incidents

Watching

Watching

1 week ago —

No more issues have been observed since the suspected faulty optic has been disabled. A technician visit to swap the optic is pending. We will continue to monitor and intervene when necessary.

Investigating

1 week ago —

We have identified an potential failed optic, showing high amount of egress drops as root cause. The interface has been disabled and LACP is currently running on degraded redundancy. As this only impacts a small subset of total network capacity, enough redundancy is still in place.

1 week ago —

We are currently observing various brief disconnects in the length of ~30 seconds each for transport towards DRT campus. To avoid further impact, several carriers have been drained. We are awaiting feedback from partner network.